OpenStudy (anonymous):

The value of y varies directly with x, and y = 12 when x = 8. Find y when x = 15.

OpenStudy (dape):

Okay so direct variation means that we can write the relation on the form \[ y = kx \] You said x=8 when y=12 that means that \[ 12 = k*8 \\ k = \frac{12}{8} = \frac{3}{2} \] So our relation is just \[ y = \frac{3}{2}x \] Putting in x=15 gives \[ y = \frac{3}{2}*15 = 22.5 \]

OpenStudy (anonymous):

22.5